Circular No. 7548

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/1963 W1 = 2000 SO_{253} (ANDERSON-LINEAR) S. Nakano, Sumoto, Japan, reports his identification of comet P/2000

SO_{253} (cf. IAUC 7524) with P/1963 W1 = 1963 IX (cf. IAUC 2013), which had been recorded on four Palomar Schmidt plates taken 1963 Nov. 22–25. The resulting orbital elements for the earlier apparition are T = 1963 Oct. 28.5 TT, q = 1.985 AU, i = 4^{o}.5, P = 7.89 yr. The comet made approaches of 0.10 and 0.40 AU from Jupiter in 1961 Aug. and 1985 Apr., respectively.

S/2000 S 12 B. Gladman and his team (cf. IAUC 7512) have discovered yet an-

other Saturnian satellite with the 3.6-m Canada-France-Hawaii telescope at Mauna Kea on Sept. 23. Complete details are on MPEC 2000-Y33.

URSID METEORS 2000 P. Jenniskens, Ames Research Center, reports that preliminary results

show that there was an enhancement of Ursid meteors visible between Dec. 22.2 and 22.4 UT, including several brighter than mag 1, with a peak (ZHR > 50) near the predicted time of Dec. 22.31 (cf. IAUC 7544).

COMETS C/2000 W2 AND C/2000 W3 (SOHO) Further to IAUC 7536, D. Hammer reports his measurements for addi-

tional Kreutz sungrazing comets found by M. Oates (C/2000 W2) and M. Meyer (C/2000 W3) on SOHO website images obtained with the C2 coronagraph. The reduced observations and orbital elements by B. G. Marsden are given on MPEC 2000-X48.

Comet  2000 UT  α_{2000}  δ_{2000}
C/2000 W2 Nov. 16.271  15^{h}25.^{m}6  −20^{o}22′
C/2000 W3  18.229  15 35.0  −20 56

COMET C/1999 T1 (McNAUGHT-HARTLEY) Visual m_{1} estimates: Dec. 8.75 UT, 8.2 (M. Mattiazzo, Wallaroo, S.

Australia, 25×100 binoculars); 16.83, 7.7 (S. Yoshida, Ibaraki, Japan, 0.25m reflector); 21.84, 7.7 (Y. Nagai, Yamanashi, Japan, 11×80 binoculars); 23.18, 7.9 (K. Hornoch, Lelekovice, Czech Republic, 10×80 binoculars).

2000 December 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green